On a map, two cities are 2 5/8 in apart.If 3/8 in on the map represents 25 mi. Find the number of miles bewteen the two cities?
This problem seems really simple but I just can't figure it out right. Hope you can help! Thank you!

Answer by mananth(16946)

It is a direct proportion sum
3/8 " = 25 miles
2 5/8 = ?
2 5/8 = 21/8 "
...
((21/8) *25)/(3/8)
(21*25/8)*8/3
(21*25*8) /8*3
=7*25
=175 miles
...
m.ananth@hotmail.ca